OpenStudy (anonymous):

f-1(3) when f(x) = 2 x minus 7, all over 3 Explain each step in solving @Hero

OpenStudy (anonymous):

@ParthKohli

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

@paki @nincompoop @Nurali

OpenStudy (nurali):

If f^-1(3) means that you have to find the inverse of f(3), than f(3) = 2x-7/3 f(3)=2*3-7/3=-1/3 and f^-1(3)=-3

OpenStudy (anonymous):

@Nurali Thank you !

OpenStudy (nurali):

My Pleasure.